OpenStudy (anonymous):

Help please... I need someone to check my homework.. In school I'm not even sure that I got it..

jimthompson5910 (jim_thompson5910):

Post what you need to have checked.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

-20=-4-6x. Which I put x=10

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

Just says ohh to wrongs. Can check it myself I just want to see see got the rig transfer an

OpenStudy (anonymous):

Okay.. I'll double checks at one

jimthompson5910 (jim_thompson5910):

-20=-4-6x -20+4=-4-6x+4 ... Add 4 to both sides. -16 = -6x -6x = -16 x = ???

OpenStudy (anonymous):

Okay.. That be...

OpenStudy (anonymous):

4 would it...

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

The teacher makes use do say an do list .. I.. To that part but I don't get the exactly rest

OpenStudy (anonymous):

Hang on let me redo this problem prob

jimthompson5910 (jim_thompson5910):

to solve -6x = -16, we divide both sides by -6 -6x = -16 -6x/(-6) = -16/(-6) x = -16/(-6) x = 8/3 ... reduce as much as possible

OpenStudy (anonymous):

Wow the teacher showed us a complete different way... Let me try it

OpenStudy (anonymous):

Go on to somewhere else I'm still trying to figure it out with my self

OpenStudy (anonymous):

Okay I will have to divide -6 from -20 that would be...

OpenStudy (anonymous):

Countining 3 so I've have to keep the division as a fraction

jimthompson5910 (jim_thompson5910):

I show that above

jimthompson5910 (jim_thompson5910):

leave it as a fraction, yes

OpenStudy (anonymous):

So then I'd have to divide (multiply) 6/20 onto 1/4 right?

jimthompson5910 (jim_thompson5910):

do you see how I went from -6x = -16 to x = 8/3 ?

OpenStudy (anonymous):

Yeah now I do

OpenStudy (anonymous):

Okay let me try finish the problem plz

OpenStudy (anonymous):

Wait wait I got 1/10

jimthompson5910 (jim_thompson5910):

the final answer is 8/3

jimthompson5910 (jim_thompson5910):

not sure how you are getting 1/10

OpenStudy (anonymous):

No I messed up hold on

OpenStudy (anonymous):

I multiplied wrong hang on in almost done

jimthompson5910 (jim_thompson5910):

the answer to -20=-4-6x is x = 8/3 or is there more to the question?

OpenStudy (anonymous):

I got stuck when I divided 80 by two... I got fourty over 3 umm I can't reduce

OpenStudy (anonymous):

Backwards 3/40

jimthompson5910 (jim_thompson5910):

what is the full question?

OpenStudy (anonymous):

Idk

jimthompson5910 (jim_thompson5910):

do you have the sheet where it gives the full question?

OpenStudy (anonymous):

That's the problem

jimthompson5910 (jim_thompson5910):

oh gotcha

jimthompson5910 (jim_thompson5910):

-20=-4x-6x

jimthompson5910 (jim_thompson5910):

the -4x and -6x combine to -10x

jimthompson5910 (jim_thompson5910):

so we have -20 = -10x

OpenStudy (anonymous):

Its the first one

jimthompson5910 (jim_thompson5910):

divide both sides by -10 to isolate x -10x = -20 x = -20/(-10) x = 2

OpenStudy (anonymous):

Well that's a coincidence because on number 2 I got -2

jimthompson5910 (jim_thompson5910):

I don't see #2

OpenStudy (anonymous):

Thanks for the help!! I'm gonna try to figure out the rest!! :)

OpenStudy (anonymous):

The teacher put the whole paper all mixed up lol thanks though!!